### v1

# FedFDP: Federated Learning with Fairness and Differential Privacy

Abstract
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, through our observations, a globally effective trained model may performance disparities in different clients. This implies that the jointly trained models by clients may lead to unfair outcomes. On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
To address the first issue mentioned above, we propose a federated algorithm with fairness, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, Extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. The code is accessible at https://anonymous.4open.science/r/FedFDP-E754.
1 Introduction
Federated Learning (FL) McMahan et al. (2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Because its ability to overcome data silos by jointly improving model performance without the need to upload local data, FL has gained significant attention in recent year Zhu et al. (2021)Lee et al. (2022). In particular, many studies have made great breakthroughs in model performance Wang et al. (2020); Li et al. (2020a, 2021b) and communication costs Wang et al. (2019); Ling et al. (2023) in FL.
However, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data Zhu et al. (2019) or the membership inference attack Song et al. (2017). Moreover, Melis et al. Melis et al. (2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in the FL, numerous scholars have begun integrating techniques such as differential privacy (DP) Dwork et al. (2014), homomorphic encryption Paillier (1999), and secure multi-party computation Yao (1982) into the federated learning framework. Among these, DP has emerged as the most widely acknowledged approach for privacy protection in federated learning, owing to its minimal computational overhead and rigorous mathematical framework. Many works Abadi et al. (2016); Carlini et al. (2019); Feldman (2020) have demonstrated that by adding proper DP noise in the training phase, the resulted machine learning models can prevent unintentional leakage of private training data, such as membership inference attacks.
Figure 1: The test accuracy of 50 clients utilizing a globally trained model in FL.
Another easily overlooked critical issue in federated learning is fairness. The globally trained model collaboratively by clients may exhibit significant performance disparities across datasets held by different clients. As shown in Fig.1, where the red dots indicate the test accuracy of the global model across various clients after the completion of federated learning using the FedAvg McMahan et al. (2017) algorithm. And we can see the globally trained model applied to individual clients, the performance significant performance disparities across differential clients. More and more scholars have begun to notice this unfair phenomenon and initiated research on it  Wang et al. (2020); Li et al. (2020b).
While numerous articles separately discuss fairness and differential privacy protection in federated learning, there is a lack of research integrating both aspects into federated learning. Balancing model performance, fairness, and differential privacy protection in federated learning presents a challenging problem. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a federated learning framework that ensures both fairness and differential privacy protection, effectively guaranteeing model performance.
Our main contributions are listed below:
  1. 1.
    Firstly, we propose a fairness-aware federated algorithm, denoted as algorithm FedFair by designing a new local loss function that is constructed to simultaneously optimize the two objectives of fairness and model performance (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Fig.1.
  2. 2.
    Building upon FedFair, we propose algorithm FedFDP to achieve differential privacy protection (Section 5) . Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. 3.
    Furthermore, we demonstrate there exist an optimal fairness parameter $\lambda$ value to balance the model performance and fairness through the convergence analysis (Section 6).
  4. 4.
    And then, we perform privacy loss analysis of our algorithm FedFDP by the notion of Rényi differential privacy (RDP) to prove that it satisfies the differential privacy (Section 7).
  5. 5.
    Lastly, through a comprehensive empirical evaluation on three public datasets, we confirm that FedFair and FedFDP outperforms the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2 Related Work
2.1 Fairness in Federated Learning
In the field of machine learning, a multitude of fairness concepts have been advanced (Mehrabi et al. Mehrabi et al. (2021), Yuxin et al. Shi et al. (2021)), each highlighting specific facets and addressing the concerns of diverse stakeholder groups. For instance, Mohri et al. Mohri et al. (2019) advocate for Good-Intent Fairness, while Zhou et al. Zhou et al. (2021) embrace Selection Fairness. Cong and Lyu et al. Cong et al. (2020)Lyu et al. (2020) employ principles from game theory Rabin (1993) to steer the allocation of rewards.
Currently, the predominant fairness criterion in Federated Learning is “Performance Parity”, as proposed by Li et al. Li et al. (2020b). This criterion aims to ensure fairness by evaluating the consistency of performance across different client devices in federated learning. This definition is widely adopted. Zhao et al. Zhao and Joshi (2022) suggested replacing the loss amplification mechanism in Li et al. (2020b) with a straightforward weight reallocation mechanism. Li et al. Li et al. (2021b) focus on optimizing the fairness of local models in personalized federated learning.
2.2 Federated Learning with Differential Privacy
Federated learning with differential privacy has become a common privacy protection method in federated learning. Enayat et al. Ullah et al. (2023) dynamically adjust the compression rate to minimize communication rounds in federated learning by responding to changes in training loss. Zhou et al. Zhou et al. (2023) investigated the effects of local iterations when utilizing the Laplace mechanism for perturbation. Fu et al. Fu et al. (2022) introduced adaptive noise adding, customizing it for different clients and communication rounds based on observed fluctuations in loss values. Ling et al. Ling et al. (2023) determined the optimal number of local epochs through convergence analysis, striking a balance between communication resources and privacy budget. Raouf et al. Kerkouche et al. (2021) improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
The aforementioned works each independently address either fairness or privacy protection in federated learning. To the best of our knowledge, we are the first to concurrently consider both fairness and privacy in the context of federated learning.
3 Preliminaries
3.1 Federated Learning
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of Federated Learning is to minimize $F(\mathbf{w})$, in other words, to seek:
$\mathbf{w}^{*}\triangleq\arg\min F(\mathbf{w}).$ (1)
The function $F(\cdot)$ is defined as follows:
$F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (2)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
3.2 DPSGD
In gradient-based methods, DPSGDAbadi et al. (2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient of loss function:
    $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$ (3)
  • Secondly, clipping the gradient of per-sample to get the sensitivity:
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right).$ (4)
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
The final step is to proceed with gradient descent as usual. Here, $C$ represents the clipping bound, and $\sigma$ is the noise multiplier. If each client in FL uses DPSGD as the gradient descent method to protect their private data, FL will transform into Differentially Private Federated Learning (DPFL).
3.3 Fairness in Federated Learning
Definition 1.
(Fairness of performance distributionLi et al. (2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally say that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (6) represents the specific form of fairness in federated learning in this paper, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients.
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
15       for $e\in[1,\cdots,E]$ do
16             for batch $b\in\mathcal{B}_{i}$ do
17                   $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18                   $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
19                   $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
20
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ and compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 1 FedFair
4 Proposed Algorithm: FedFair
In the previous section, we identified the performance objective function as shown in Equation (2) and the fairness objective function as shown in Equation (6). Both of these objective functions are composed of individual client objectives. In this section, we will focus on the performance and fairness of each client and combine these two objectives to construct a new objective function, based on which we propose a fair federated learning algorithm.
We simultaneously optimize performance and fairness, and can get the optimization objective for client $i$ as:
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}_{t})\right)^{2},$ (7)
where $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t-1})$ [^1] The reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.and $\lambda$ is a hyperparameter used to adjust the degree of fairness, $\lambda\geq 0$. When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
The gradient of Equation (7) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t})$.
The stochastic gradient descent for Equation (7) as show:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Algorithm 1.
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For $E$ epochs and each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
Furthermore, as the iterations advance, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods Zinkevich (2003).
Figure 2: FedFDP framework
5 Proposed Algorithm: FedFDP
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Fig. 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows for adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm Fu et al. (2023); Abadi et al. (2016); Bu et al. (2022); Yang et al. (2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})\right)^{2},$ (10)
where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
Next, we compute the gradient of Equation (10):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{loss}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
$=\left(1+\lambda\cdot\Delta_{loss}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
where $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$.
To achieve differential privacy protection, we replace loss function gradient by Equation (5.1) based on Equations (3)-(5), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta^{j}_{loss}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments 5.
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{loss}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $C_{l}^{i,t}\cdot\sigma_{l}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t})$ is challenging. A large $C_{l}^{i,t}$ helps avoid loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
where
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))),$ (15)
To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$. Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
5.3 Overall Algorithm
We utilize the equation form presented in (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for training $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}=\text{Initial()}$();
2SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
16             $\Delta_{loss}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
17             Compute $C^{i,j}_{t}$ by Equation (13);
18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
19
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\mathcal{N}(0,\mathbf{I})\right)$;
21       Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
23             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
24             $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
25
26      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
27       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
28
Algorithm 2 FedFDP
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6 Utility Analysis
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal $\lambda$ that minimizes the convergence bound.
6.1 Convergence Analysis
We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound. Prior to proving this, we need to establish several assumptions:
Assumption 1.
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2.
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3.
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=1,\cdots,N$ , $t=1,\cdots,T-1$ and $j=1,\cdots,|\mathcal{B}_{i}|$
Theorem 1.
Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
The meaning of $\Gamma$ is consistent with Li et al. (2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (16). For the detailed proof, please refer to Appendix A.1.
6.2 Optimal $\lambda$
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{loss}^{j}$. We assume $Q_{0}\leq\Delta_{loss}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in (16) as a function of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{2Q_{0}\lambda+1},$ (17)
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$.
We sketched a rough graph of $P(\lambda)$, where $\lambda\in[0,+\infty)$. Upon observing Fig. 3, it becomes evident that when $a_{1}$,$a_{2}$,$a_{3}$,$a_{4}$ greater than 00, there exists an optimal value of $\lambda$ that minimizes the right-hand side of Equation (16), aligning with the experimental results presented in Fig. 4.
Figure 3: An example of $P(\lambda)$
7 Privacy Analysis
Because each client’s locally trained pair of data samples has the same sampling rate and noise multiplier in each round, each client has the same privacy loss in each round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 2.
(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3 in Appendix A.2, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4 in Appendix A.2.
The details of the proof please refer to Appendix A.2.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
Table 1: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 79.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
8 Experiments
In this section, we conduct experiments to demonstrate the performance and fairness of FedFair and FedFDP using three real datasets. Additionally, we investigate the impact of various factors on FedFDP, such as fairness hyperparameter $\lambda$, original clipping bound $C$, heterogeneity and scalability. The impact of heterogeneity and scalability please refer to Appendix A.3).
8.1 Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST LeCun et al. (1998), FashionMNIST Xiao et al. (2017), and CIFAR10 Krizhevsky (2009). And we comparised our algorithms to six baselines, which include FedAvgMcMahan et al. (2017), SCAFFLODKarimireddy et al. (2020), FedProxLi et al. (2020a), FedDynAcar et al. (2021), ALI-DPFLLing et al. (2023) and q-FFLLi et al. (2020b). For the details of parameters settings and the descriptions of baselines please refer to Appendix A.3.1.
8.2 Results
Performance and Fairness. The experimental results of the FedFair algorithm and baseline methods are shown in Table 1. FedFair notably enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.5%. And Table 2 shows the experimental results of the FedFDP algorithm and baseline methods. Under the same privacy budget, FedFDP demonstrates improvements in model accuracy of 1.2%, 1.8%, and 0.4% compared to the best baseline on the three datasets, while also enhancing fairness by 30.3%, 6.7%, and 27.8% respectively compared to the best baseline.
Figure 4: Impact of different $\lambda$ in FedFDF algorithm
Impact of different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in FedFDF algorithm. As depicted in Figure 4, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase.
Figure 5: Impact of different $C$ in FedFDF algorithm
Impact of different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 5, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values of [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, at this point, due to excessive noise addition, the model does not converge. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good values.
9 Conclusion
In this paper, firstly, we propose a FedFair algorithm effectively addresses fairness concerns by optimizing a novel local loss function. Building upon FedFair, we propose the FedFDP, incorporates a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values to achieve both fairness and differential privacy protection. And then, through convergence analysis, we determine the existence of an optimal fairness parameter, $\lambda$, striking a balance between model performance and fairness. We conduct a comprehensive privacy analysis of our approach using RDP and validate our algorithms through extensive experiments. The results indicate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. We believe our work contributes a valuable FL framework for addressing fairness and privacy challenges.
Appendix A Appendix
In the beginning, Table 3 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
Table 3: Summary of main notations
$F(\mathbf{w}$)
Global loss function
$F_{i}(\mathbf{w}$)
Local loss function for client $i$
$T$
Communication rounds
$t$
Communication round index
$D_{i}$
Dataset of client $i$
$D$
The union of $D_{i}$
$\mathbf{w}_{t}^{i}$
Local model parameter at node $i$ at iteration $t$
$\mathbf{w}^{*}$
Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$
Gradient descent step size
$\lambda$
Fairness parameter
$N$
Total number of client
$q$
batch sample ratio
$|\mathcal{B}_{i}|$
Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$d$
The dimension of model
$p_{i}$
Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$
Single data sample in $D_{i}$, indexed by $j$
$\sigma$
Noise multiplier for gradient
$\sigma_{l}$
Noise multiplier for loss
$C$
Original clipping bound for gradient
$C_{l}$
Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | |
$L_{2}$-norm
A.1 Proof of Theorem 1
For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$. We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step. Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (12) to:
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (19)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta^{j}_{loss},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by Stich [2018]. Therefore,
$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (20)
A.1.1 Key Lemma
Lemma 1.
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 1, this implies:
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
$\leq\frac{t-1}{t^{2}}v$
$\leq\frac{v}{t+1}$
Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
A.1.2 Proof of Lemma 1
By the Equation (A.1), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$*\displaystyle=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
Firstly, we process $\mathcal{A}_{2}$:
$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
Since $\mathbb{E}\mathcal{B}_{0}=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}\mathcal{B}_{2}\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
taking exception and according to assumption 3:
$\mathbb{E}\mathcal{B}_{1}\leq\eta^{2}\mathbb{E}\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\leq\eta^{2}C_{t}^{2}G^{2},$
where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
$\mathbb{E}\mathcal{A}_{2}$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}\mathcal{C}_{0}=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
$\mathbb{E}\mathcal{C}_{1}\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
According to Assumption 2, we know that
$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
Next, we proceed to handle $\mathcal{D}_{1}$.
$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
where L-smooth inference as show:
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (21)
Thus, we get
$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
To sum up,
$\mathbb{E}\mathcal{A}_{1}$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
$\quad\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}$
$=\mathbb{E}\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}$
$\leq\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}$
$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
$=\mathbb{E}\mathcal{A}_{2}=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
So we get,
$\mathbb{E}\mathcal{A}_{1}$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
Table 4: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.67 |||| 4.1e9 75.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 72.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 71.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 73.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 77.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 82.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 95.13 |||| 2.3e10 94.56 |||| 3.4e9 83.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
Table 5: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
A.2 Proof of Theorem 2
As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
A.2.1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 3.
After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (22)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 2.
(RDP privacy budget of SGMMironov et al. [2019]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (23)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (24)
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work Mironov et al. [2019] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (25)
Definition 3.
(Composition of RDPMironov [2017]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 2 and Definition 3, the Theorem 3 is obtained.
A.2.2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 4.
After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (26)
where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
A.2.3 Proof Loss of FedFDP
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 2 is hold.
Lemma 2.
(Conversion from RDP to DP Balle et al. [2020]). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
A.3 Experiment Settings
A.3.1 Parameters Settings
For all datasets, we adopted a widely used heterogeneous settingsLin et al. [2020]Li et al. [2021a] to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, where a smaller $\beta$ indicates a higher degree of heterogeneity. We set the default value of $\beta=0.1$ Lin et al. [2020]Wang et al. [2020].
In all datasets, we use a 4-layer CNN architecture McMahan et al. [2017] which consists of two convolutional layers and two fully connected layers. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
In all experiments, we use 25% of the data will form the test dataset, and the remaining 75% of the data will be used for training. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
A.3.2 Baselines introduction
The baseline methods included in the comparison are:
  • FedAvgMcMahan et al. [2017]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFDP into the FedAvg algorithm.
  • SCAFFLODKarimireddy et al. [2020]: SCAFFOLD uses control variates (variance reduction) to correct for the ’client-drift’ in its local updates. It requires significantly fewer communication rounds and is not affected by data heterogeneity or client sampling.
  • FedProxLi et al. [2020a]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  • FedDynAcar et al. [2021]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  • ALI-DPFLLing et al. [2023]: ALI-DPFL dynamically obtains the optimal number of epochs for each round of communication by minimizing the convergence upper bound, thereby addressing the imbalance between local privacy budgets and communication resources.
  • q-FFLLi et al. [2020b]: q-FFL amplifies the loss values of clients, equivalent to giving larger update weights to poorly performing clients, in order to cater to the underperforming clients, uniforming model performance across multiple devices.
A.3.3 Herterogeneity and Scalability
As shown in Table 4 and 5, we conducted additional experiments on the MNIST, FashionMNIST, and CIFAR10 datasets to compare the performance of FedFDP with that of six other algorithms in terms of Heterogeneity and Scalability. The Heterogeneity experiments were mainly conducted in three heterogeneous scenarios, extending from the default $Dir(0.1)$ to include $Dir(0.5)$ and $Dir(1)$. Scalability was expanded from the default 10 clients to include 20 and 50 clients.
Heterogeneities. Tables 4 and 5 show that when confronted with diverse data distributions exhibiting varying levels of heterogeneity, both FedFair and FedFDP outperform baseline methods in accuracy while maintaining significant fairness. With the exception of the $Dir(0.5)$ scenario, where they fall slightly short of q-FFL, they respectively demonstrated fairness improvements of 17.0% and 50% for FedFair and 30.3% and 18.9% for FedFDP compared to the best baselines in $Dir(0.1)$ and $Dir(1.0)$.
Scalability. Table 4 demonstrates that, with accuracy mostly superior to baselines, FedFair exhibited fairness improvements of 9.5%, 15.4%, and 0.01% across three client counts. Table 5 indicates that as the number of clients increases, FedFDP outperformed the best baselines by 0.36%, 0.18%, and 0.59% in accuracy, except in the 50 clients scenario, where it demonstrated fairness improvements of 27.8% and 24.1%.