### v2

# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

Abstract.
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, through our observations, a globally effective trained model may performance disparities in different clients. This implies that the jointly trained models by clients may lead to unfair outcomes. On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
To address the first issue mentioned above, we propose a fairness-aware federated learning algorithm, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Code and data is accessible at https://anonymous.4open.science/r/FedFDP-5607.
Federated Learning, Differential Privacy, Fairness
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Federated Learning (FL) (McMahan et al., 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Thanks to its ability to overcome data silos by jointly improving model performance without the need to upload raw data, FL has gained significant attention in recent years (Zhu et al., 2021)(Lee et al., 2022). In particular, many studies have made great breakthroughs in model performance (Wang et al., 2020; Li et al., 2020b, 2021b) and communication costs (Wang et al., 2019; Ling et al., 2023) in FL.
Figure 1. The test accuracies of 50 clients utilizing a globally trained model in FL.
[…]
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al., 2021c). Current machine learning algorithms that protect fairness, such as  (Dwork et al., 2012; Cotter et al., 2019), have not adequately addressed the issue of performance disparities among clients in federated learning. As shown in Figure 1, the red dots represent the test accuracies of the global model on different clients after completing federated learning with the FedAvg algorithm (McMahan et al., 2017). The global model trained collaboratively by clients exhibits significant unbalanced performance across different datasets held by the various clients. An increasing number of scholars are beginning to pay attention to this performance disparities phenomenon and are starting to work on it  (Li et al., 2020c, 2021b; Zhao and Joshi, 2022).
On the other hand, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data (Zhu et al., 2019) or the membership inference attack (Song et al., 2017). Moreover, Melis et al. (Melis et al., 2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in FL, numerous scholars have begun integrating techniques such as differential privacy (DP) (Dwork et al., 2014), homomorphic encryption (Paillier, 1999), and secure multi-party computation (Yao, 1982) into the federated learning framework. Among these, DP stands out as the premier approach to privacy preservation in federated learning, attributed to its modest computational overhead and robust mathematical underpinnings. This is particularly relevant in defending membership inference attacks, as several studies (Abadi et al., 2016; Carlini et al., 2019; Feldman, 2020) have shown that incorporating appropriate DP during the training phase can effectively prevent inadvertent disclosure of sensitive training information.
Although numerous articles have addressed fairness and differential privacy individually within the context of federated learning, there is a scarcity of research that integrates these two aspects. Achieving a balance between model performance, fairness, and differential privacy protection within federated learning frameworks poses a significant challenge. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a fairness-aware federated learning with differential privacy, effectively guaranteeing model performance when achieving fairness and privacy protection.
Our main contributions are listed as follows:
  1. (1)
    Firstly, a fairness-aware federated algorithm is proposed, denoted as FedFair. It incorporates a novel loss function, designed to optimize the objectives of fairness and model performance simultaneously (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Figure1.
  2. (2)
    Building upon FedFair, an algorithm referred to as FedFDP is proposed to achieve differential privacy protection (Section 5). Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. (3)
    Furthermore, it is demonstrated that an optimal fairness parameter $\lambda$ is found, which, when determined, enables the fastest convergence of FedFDP in balancing model performance and fairness through the convergence analysis (Section 6).
  4. (4)
    And then, a privacy loss analysis is conducted on algorithm FedFDP using the notion of Rényi differential privacy (RDP) to verify its satisfaction of differential privacy requirements (Section 7).
  5. (5)
    Lastly, through a comprehensive empirical evaluation on three public datasets, it is confirmed that FedFair and FedFDP outperform the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2. Related Work
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Ullah et al., 2023; Zhou et al., 2023; Fu et al., 2022; Ling et al., 2023; Kerkouche et al., 2021; Fu et al., 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri et al., 2019) and Selection Fairness (Zhou et al., 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al., 2020; Lyu et al., 2020; Mehrabi et al., 2021; Shi et al., 2021).
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al., 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al., 2021; Padala et al., 2021; Gu et al., 2022, 2020; Pentyala et al., 2022).
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al., 2020c). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi, 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al., 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
In comparison to definitions such as group fairness, balanced performance fairness is more meaningful for data providers such as enterprises and institutions, moreover, it can effectively motivate clients participate in the federated system. However, achieving balanced performance fairness typically requires adjusting the direction of model updates. It is a challenge to refine the model for balanced performance fairness while integrating DP techniques, such as clipping and noise addition, into the model’s update process. To the best of our knowledge, we are the first to consider the work of balanced performance fairness in DPFL.
3. Preliminaries
In the beginning, Table 1 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
Table 1. Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
3.1. Federated Learning and Fairness
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
(1) $\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
We have defined balanced performance fairness in FL as follows. Specifically, this type of fairness is defined based on the performance variance of the model across clients: the smaller the variance, the fairer the model.
Definition 0.
(Balanced performance fairness (Li et al., 2020c)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
3.2. Differential Privacy
Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
Definition 0.
(Differential Privacy (Dwork et al., 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
(2) $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow. We use $l_{2}$ sensitivity as in the previous work.
Definition 0.
($\mathbf{l_{k}}$-Sensitivity(Dwork et al., 2006)) For a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, we define its $l_{k}$ norm sensitivity (denoted as $\Delta_{k}f$) over all neighboring datasets $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ differing in a single sample as
(3) $\text{sup}_{D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}}||f(D_{i})-f(D_{i}^{\prime})||_{k}\leq\Delta_{k}f.$
DPSGD. In gradient-based methods, DPSGD(Abadi et al., 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  1. (1)
    Firstly, computing the gradient of loss function:
    (4) $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$
  2. (2)
    Secondly, clipping the gradient of per-sample to get the sensitivity:
    (5) $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$
    where $C$ represents the clipping bound.
  3. (3)
    Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    (6) $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$
The final step is to proceed with gradient descent as usual. When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
20
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 1 FedFair
4. Proposed Algorithm: FedFair
We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1. Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
In this study, we employ the loss of model denote the performance. Equation (7) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
(7) $\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
Upon incorporating Equation (7) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
(8) $\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
For client $i$ during round $t$, the refined objective is articulated as follows:
(9) $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in section 6.2. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (9), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (9) as follows:
(10) $\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (10) as show:
(11) $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Figure 2. FedFDP framework.
[…]
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Algorithm 1.
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich, 2003).
5. Proposed Algorithm: FedFDP
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Figure 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
17             Compute $C^{i,j}_{t}$ by Equation (15);
18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
19
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
21
22      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (16) and (17) ;
23       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
24             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
25
26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
27
28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\cdot\mathcal{N}(0,\mathbf{I})\right)$;
29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
30
Algorithm 2 FedFDP
5.1. Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al., 2016; Fu et al., 2023; Bu et al., 2022; Yang et al., 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (9), into Equation (12):
(12) $\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$
where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
Next, we compute the gradient of Equation (12):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{i}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
(13) $=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To achieve differential privacy protection, we replace the gradient of loss function by Equation (5.1) based on Equations (4)-(6), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
(14) $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
where:
(15) $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (15) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (15) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (15) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments (Figure 4).
5.2. Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
(16) $C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$
where
(17) $\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$
To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$. Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
5.3. Overall Algorithm
We utilize the equation form Equation (14) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6. Optimal Fairness Parameter
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal fairness parameter $\lambda$ that minimizes the convergence bound.
6.1. Convergence Analysis
We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound. Prior to proving this, we need to establish several assumptions:
Assumption 1.
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2.
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3.
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
Theorem 1.
Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
(18) $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
The meaning of $\Gamma$ is consistent with (Li et al., 2020a), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (18). For the detailed proof, please refer to Appendix A.1333Due to the lengthy proof, please refer to https://anonymous.4open.science/r/FedFDP-5607/APPENDIX/FedFDP_appendix.pdf for the appendix..
6.2. Optimal $\lambda$
Theorem 2.
There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 2.
Proof. When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (18) to obtain a function in terms of $\lambda$,
(19) $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term. Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized. The derivative of $\mathcal{F}(\lambda)$ is as follows:
$\mathcal{F}^{\prime}(\lambda)$ $=\frac{(3a_{1}\lambda^{2}+2a_{2}\lambda+a_{3})(a_{5}\lambda+1)}{(a_{5}\lambda+1)^{2}}-\frac{a_{5}(a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4})}{(a_{5}\lambda+1)^{2}}$
(20) $=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
(21) $\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$
Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
(22) $=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$
The derivative of $\mathcal{G}(\lambda)$ is given by
(23) $\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
(24) $\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$
where
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
[…]
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
7. Privacy Analysis
As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
7.1. Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 1.
After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
(25) $R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$
where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
The detail of proof please refer Appendix A.2.
7.2. Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 2.
After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
(26) $R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$
where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
7.3. Privacy Loss of FedFDP
Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 4 to convert it to $(\epsilon,\delta)$-DP. Lastly, we can get the Privacy loss of FedFDP as follows.
Theorem 3.
(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 2.
Lemma 0.
(Conversion from RDP to DP (Balle et al., 2020)). if a randomized mechanism $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ satisfies $(\alpha,R)$-RDP , then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
8. Experiments
Table 2. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
Table 3. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
Table 4. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
Table 5. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
[…]
$\epsilon$ 1.0 2.0 3.0 4.0
Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
This section outlines our experimental framework, beginning with the presentation of baselines and default parameter configurations. We then detail the experimental assessment of FedFair and FedFDP across dimensions of performance, fairness, heterogeneity, and scalability. Ablation studies focusing on the hyperparameters $\lambda$, $C$ and $\epsilon$ of FedFDP are also included. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
8.1. Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009). And we comparised our algorithms to six baselines:
  1. (1)
    FedAvg(McMahan et al., 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. (2)
    SCAFFLOD(Karimireddy et al., 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. (3)
    FedProx(Li et al., 2020b): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. (4)
    FedDyn(Acar et al., 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. (5)
    ALI-DPFL(Ling et al., 2023): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. (6)
    q-FFL(Li et al., 2020c): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al., 2020; Li et al., 2021a) to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, and the default value of $\beta=0.1$ (Lin et al., 2020; Wang et al., 2020). We use a 4-layer CNN architecture (McMahan et al., 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
8.2. Results
In this subsection, we systematically evaluate the performance and fairness of the FedFair and FedFDP under varying heterogeneity levels, expanding from $Dir(0.1)$ to both $Dir(0.5)$ and $Dir(1)$. We also assess their scalability by scaling the number of clients from the standard 10 to encompass 20 and 50 clients. Furthermore, we examine the influence of three pivotal hyperparameters within the FedFDP algorithm, namely, the adjustments to $\lambda$, $C$ and $\epsilon$.
8.2.1. FedFair
Performance and Fairness. Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
Scalability. Table 4 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
8.2.2. FedFDP
Performance and Fairness. Table 3 displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table 5 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
Scalability. Table 5 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Figure 3. Impact of different $\lambda$ in FedFDP algorithm
[…]
Figure 4. Impact of different $C$ in FedFDP algorithm
[…]
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and fairness $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis in Section 6.1: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. In Section 5.1, we emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 6. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
9. Conclusion
In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function. The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection. Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness. Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP. Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness. It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
Appendix A Appendix
A.1. Proof of Theorem 1
For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$. We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step. Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (14) to:
(28) $\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich, 2018). Therefore,
$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
(29) $\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
A.1.1. Key Lemma
Lemma 0.
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 1, this implies:
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
$\leq\frac{t-1}{t^{2}}v$
$\leq\frac{v}{t+1}$
Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
A.1.2. Proof of Lemma 1
By the Equation (A.1), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
Firstly, we process $\mathcal{A}_{2}$:
$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
taking exception and according to assumption 3:
$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
According to Assumption 2, we know that
$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
Next, we proceed to handle $\mathcal{D}_{1}$.
$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
where L-smooth inference as show:
(30) $\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$
Thus, we get
$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
To sum up,
$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
So we get,
$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
A.2. Proof of Theorem 1
$proof.$ We will prove Theorem 1 in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 0.
(RDP privacy budget of SGM(Mironov et al., 2019)). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
(31) $R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$
where
(32) $\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work (Mironov et al., 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
(33) $A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$
Definition 0.
(Composition of RDP(Mironov, 2017)). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 2 and Definition 3, the Theorem 1 is obtained.