### v3

# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

Abstract
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, federated learning faces challenges in fairness and data privacy. To address both of the above challenges simultaneously, we first propose a fairness-aware federated learning algorithm, termed FedFair. Then based on FedFair, we introduce differential privacy protection to form the FedFDP algorithm to address the trade-offs among fairness, privacy protection, and model performance. In FedFDP, we designed an fairness-aware gradient clipping technique to identify the relationship between fairness and differential privacy. Through convergence analysis, we determined the optimal fairness adjustment parameters to simultaneously achieve the best model performance and fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping method to minimize privacy budget consumption. Extensive experimental results demonstrate that FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Codes and datasets will be made public after acceptance.
1 Introduction
Federated Learning (FL) (McMahan et al. 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets. After training their local models, each client only transmits model parameters to the central server, rather than the original training data. The server aggregates these parameters to update the global model, which is then sent back to the clients for further training. Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years (Zhu, Hong, and Zhou 2021; Lee et al. 2022; Chen et al. 2024). Particularly, many studies have achieved significant breakthroughs in terms of model performance (Wang et al. 2020; Li et al. 2020a, 2021b) and communication costs (Wang et al. 2019; Ling et al. 2024) in FL.
Figure 1: Client drift caused by Non-IID.
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al. 2021a). Current fairness-aware machine learning algorithms, such as (Dwork et al. 2012; Cotter et al. 2019), have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (yellow squares) lead to significant performance imbalances of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity (Li et al. 2020b, 2021b; Zhao and Joshi 2022).
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data (Zhu, Liu, and Han 2019) or member inference attacks (Song, Ristenpart, and Shmatikov 2017). Even if only models or gradients are uploaded instead of original data, data leakage can still occur (Melis et al. 2022). To enhance data privacy in FL, differential privacy (DP) (Dwork, Roth et al. 2014) has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information (Abadi et al. 2016; Carlini et al. 2019; Feldman 2020).
Although some works have explored fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework. For the first time, we propose a fairness-aware federated learning with differential privacy. However, balancing differential privacy protection, fairness, and model performance presents a significant challenge. Specifically, there are two main challenges: First, the clipping and noise addition steps in differential privacy can impact fairness and model utility in federated learning. Second, fair federated learning requires uploading loss information to the central server, which incurs additional privacy budget consumption.
In this paper, to address the first challenge, we analyze the relationship between the fairness loss function and differential privacy in federated learning. We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy. This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy. Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness. To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
Our main contributions are as follows:
  • First, we propose a fairness-aware federated algorithm named FedFair. It combines a novel loss function aimed at optimizing the objectives of fairness and model performance simultaneously.
  • Based on FedFair, we propose an algorithm called FedFDP to further equip the system with differential privacy. In particular, we designed a fairness-aware gradient clipping strategy that ensures differential privacy and allows for adjustment of fairness in federated learning. In addition, we propose an adaptive clipping method for the additional loss values uploaded by each client to achieve optimal utility.
  • Further, we conducted a convergence analysis of FedFDP and found the optimal fairness parameter $\lambda$ by minimizing the upper bound of convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
  • Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness. What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
2 Related Work
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Fu, Chen, and Han 2022; Zhou et al. 2023; Ling et al. 2024; Fu et al. 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri, Sivek, and Suresh 2019) and Selection Fairness (Zhou, Fang, and Hui 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al. 2020; Lyu et al. 2020) .
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al. 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al. 2021; Padala, Damle, and Gujar 2021; Gu et al. 2022, 2020; Pentyala et al. 2022).
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al. 2020b). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al. 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
3 Preliminaries
3.1 Federated Learning and Fairness
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
We have defined balanced performance fairness in FL as follows.
Definition 1
(Balanced performance fairness (Li et al. 2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
3.2 Differential Privacy
Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
Definition 2
(Differential Privacy (Dwork, Roth et al. 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow. We use $l_{2}$ sensitivity as in the previous work.
DPSGD. In gradient-based methods, DPSGD(Abadi et al. 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient of loss function: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
  • Secondly, clipping the gradient of per-sample to get the sensitivity: $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right)$, where $C$ represents the clipping bound.
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier: $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right)$.
The final step is to proceed with gradient descent as usual. When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
4 FedFair: Equipping FL with Fairness
We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1. Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
Upon incorporating Equation (2) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (4)
For client $i$ during round $t$, the refined objective is articulated as follows:
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (5)
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 2 of Algorithm 2, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (5), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (5) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (6)
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (6) as show:
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (7)
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Figure 2: FedFDP framework.
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Appendix Algorithm 2.
5 FedFDP: Further Equipping FedFair with Differential Privacy
In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy. The flow of FedFDP algorithm is shown in Figure 2. It adds two additional processes to FedFair:
  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
Next, we discuss the details of these two strategies.
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
17             Compute $C^{i,j}_{t}$ by Equation (11);
18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
19
20      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
21
22      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (12) and (13) ;
23       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
24             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
25
26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
27
28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
30
Algorithm 1 FedFDP
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al. 2016; Fu et al. 2023; Bu et al. 2022; Yang et al. 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (5), into Equation (8):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (8)
where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
Next, we compute the gradient of Equation (8):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (9)
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To ensure the algorithm is differentially private, we replace the gradient of loss function by Equation (9), and reformulate the computation of the gradient descent for client $i$ as follows:
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (10)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (11)
Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (11) corresponds to the traditional DP clipping method without fairness. For large $C$, Equation (11) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (11) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 1-1 of Algorithm 1, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$. For the $i$-th client, the adaptive clipping method uses the $differentially$ $private$ $mean$ $of$ $the$ $previous$ $round$ as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (12)
where
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (13)
To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$. Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
5.3 Overall Algorithm
We utilize the equation form Equation (10) to enhance Algorithm 1. Compared with the Algorithm 2, FedFDP solely necessitates the application of differential privacy to gradient and loss.
  • DPSGD (lines 1-1): In lines 1-1, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 1 and execute the gradient decent.
  • DP for local loss (lines 1-1): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6 Finding Optimal Fairness Parameter
In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
Table 1: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
6.1 Convergence Analysis
We analyze the convergence of Algorithm 1 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound. Prior to proving this, we need to establish several assumptions:
Assumption 1
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
Theorem 1
Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 1 as follows:
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (14)
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
The meaning of $\Gamma$ is consistent with (Li et al. 2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (14). For the detailed proof, please refer to Appendix.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
$\epsilon$ 1.0 2.0 3.0 4.0
Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
[…]
6.2 Optimal Fairness Parameter $\lambda$
Based on the convergence analysis, we find the “best” fairness parameter $\lambda$ that can minimize the convergence bound. We have the following theorem to do so:
Theorem 2
There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 1.
proof sketch: By reorganizing the convergence upper bound of Equation (14) into a function of $\lambda$:
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (15)
By using numerical analysis methods, the minimum value $P(\lambda)$ can be solved. For the meaning of $a_{1,2,3,4,5}$ and the proof process, please refer to the Appendix.
7 Privacy Analysis
Because each client has the same privacy loss in every round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 3
(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 4 in Appendix, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 5 in Appendix.
The details of the proof please refer to Appendix.
8 Experiments
This section outlines our experimental framework. We detail the experimental assessment of FedFDP across dimensions of performance, fairness, heterogeneity, and scalability, including ablation experiments focusing on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$ and noise multiplier $\sigma$.
In the appendix, we also provide an experimental assessment of FedFair across the same dimensions of performance, fairness, heterogeneity, and scalability. For an introduction to baselines, default values of hyperparameters, and the running environment, please refer to the Appendix.
All tasks were repeated five times and report the mean and standard deviation.
8.1 Results
Performance and Fairness. Table 1 displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table 2 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table 2 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
Figure 4: Impact of different $C$ in FedFDP algorithm.
Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 3. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
Impact of Different $\sigma$. We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication. As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence. When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value. As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
9 Conclusion
In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function. The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection. Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness. Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP. Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness. It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
Appendix A Appendix
In the beginning, Table 4 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
Table 4: Summary of main notations
A.1 Extra Experiment
We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
Table 5: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
FedAvg 98.67 |||| 7.3e-3 98.37 |||| 1.2e-3 89.41 |||| 3.6e-2 62.03 |||| 1.1e0 60.89 |||| 1.5e0 59.78 |||| 1.5e0
SCAFFLOD 98.45 |||| 5.0e-3 98.51 |||| 2.1e-3 86.39 |||| 3.4e-2 62.21 |||| 9.9e0 59.48 |||| 8.7e0 58.64 |||| 2.3e0
FedProx 98.70 |||| 7.5-3 98.35 |||| 1.2e-3 88.33 |||| 3.5e-2 62.13 |||| 1.2e0 60.33 |||| 2.2e0 61.31 |||| 1.6e0
FedDyn 98.40 |||| 8.0e-3 98.36 |||| 1.8e-3 88.26 |||| 3.5e-2 62.33 |||| 9.4e-1 61.35 |||| 2.6e0 62.52 |||| 1.5e0
q-FFL 98.72 |||| 4.7e-3 98.61 |||| 1.1e-3 89.03 |||| 3.6e-2 63.33 |||| 9.4e-1 60.63 |||| 1.3e0 59.65 |||| 1.5e0
FedFair 98.75 |||| 3.9e-3 98.72 |||| 1.2e-3 89.60 |||| 1.7e-2 62.38 |||| 8.5e-1 61.58 |||| 1.1e0 62.90 |||| 1.5e0
Table 6: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al. 1998), FashionMNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR10 (Krizhevsky 2009). And we comparised our algorithms to six baselines:
  1. 1.
    FedAvg(McMahan et al. 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. 2.
    SCAFFLOD(Karimireddy et al. 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. 3.
    FedProx(Li et al. 2020a): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. 4.
    FedDyn(Acar et al. 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. 5.
    ALI-DPFL(Ling et al. 2024): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. 6.
    q-FFL(Li et al. 2020b): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al. 2020; Li, He, and Song 2021) to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ (Lin et al. 2020; Wang et al. 2020). We use a 4-layer CNN architecture (McMahan et al. 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
In hyper-parameters, for FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
Extended description of FedFair
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
20
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 2 FedFair
  • Server (lines 2-2): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 2-2): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 2-2. For line 2, we execute gradient descent by using $\eta_{i}$.
Result of FedFair
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich 2003). Performance and Fairness. Table 5 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 6 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table 6 indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
A.2 Proof of Theorem 1
For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$. We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step. Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (10) to:
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (17)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by (Stich 2018). Therefore,
$\mathbf{v}_{t+1}$ $=\mathbf{w}_{t}$
$\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (18)
Key Lemma
Lemma 1
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 1, this implies:
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}$ $\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}$
$=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)$
$\leq\frac{t-1}{t^{2}}v$
$\leq\frac{v}{t+1}$
Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
Proof of Lemma 1
By the Equation (A.2), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$=\|\mathbf{w}_{t}-\mathbf{w}^{*}$
$-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
Firstly, we process $\mathcal{A}_{2}$:
$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
taking exception and according to assumption 3:
$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\quad\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}$
$\quad\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
According to Assumption 2, we know that
$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot$
$\quad\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)$
$\quad-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
Next, we proceed to handle $\mathcal{D}_{1}$.
$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]$
$\quad+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
where L-smooth inference as show:
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (19)
Thus, we get
$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
To sum up,
$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
So we get,
$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}$
$\quad+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
A.3 Proof of Theorem 2
When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (14) to obtain a function in terms of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (20)
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term. Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized. The derivative of $\mathcal{F}(\lambda)$ is as follows:
$\mathcal{F}^{\prime}(\lambda)=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$ (21)
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
$\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$ (22)
Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
$=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$ (23)
The derivative of $\mathcal{G}(\lambda)$ is given by
$\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$ (24)
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
$\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$ (25)
where
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
[…]
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{1,2,3}=x_{1,2,3}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
A.4 Proof of Theorem 3
As Algorithm 1 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 1-1 and lines 1-1 of the Algorithm 1, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
Privacy Loss of $\mathbf{w}_{t+1}^{i}$.
Theorem 4
After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (26)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 3
(RDP privacy budget of SGM(Mironov, Talwar, and Zhang 2019)). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (27)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (28)
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}\text{ satisfies }\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work (Mironov, Talwar, and Zhang 2019) describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (29)
Definition 4
(Composition of RDP(Mironov 2017)). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 3 and Definition 4, the Theorem 4 is obtained.
Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
Theorem 5
After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (30)
where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
Proof Loss of FedFDP.
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 2 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 3 is hold.
Lemma 2
(Conversion from RDP to DP (Balle et al. 2020)). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.