### v4

# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

Abstract
Federated learning (FL) is an emerging machine learning paradigm designed to address the challenge of data silos, attracting considerable attention. However, FL encounters persistent issues related to fairness and data privacy. To tackle these challenges simultaneously, we propose a fairness-aware federated learning algorithm called FedFair. Building on FedFair, we introduce differential privacy to create the FedFDP algorithm, which addresses trade-offs among fairness, privacy protection, and model performance. In FedFDP, we developed a fairness-aware gradient clipping technique to explore the relationship between fairness and differential privacy. Through convergence analysis, we identified the optimal fairness adjustment parameters to achieve both maximum model performance and fairness. Additionally, we present an adaptive clipping method for uploaded loss values to reduce privacy budget consumption. Extensive experimental results show that FedFDP significantly surpasses state-of-the-art solutions in both model performance and fairness.
Index Terms:
Federated Learning, Differential Privacy, Fairness
I Introduction
Federated Learning (FL) [33] is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets. After training their local models, each client only transmits model parameters to the central server, rather than the original training data. The server aggregates these parameters to update the global model, which is then sent back to the clients for further training. Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years [54, 23, 6]. Particularly, many studies have achieved significant breakthroughs in terms of model performance [46, 27, 26] and communication costs [47, 31] in FL.
Figure 1: Client drift caused by Non-IID.
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [25]. Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [28, 26, 50].
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data [53] or member inference attacks [42]. Even if only models or gradients are uploaded instead of original data, data leakage can still occur [35]. To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [1, 5, 12].
Although some works have explored balanced performance fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework. For the first time, we propose a fairness-aware federated learning with differential privacy. However, balancing differential privacy protection, fairness, and model performance presents a significant challenge. Specifically, there are two main challenges: First, the clipping and noise addition processes in differential privacy have the potential to affect both fairness and model utility in federated learning. Moreover, achieving fairness in federated learning requires uploading loss information to the server, which entails additional privacy budget consumption.
In this paper, the relationship between the fairness loss function and differential privacy in federated learning is analyzed to address the first challenge. We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy. This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy. Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness. To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
Our main contributions are as follows:
  1. 1.
    First, a fairness-aware federated algorithm, FedFair, is proposed. It integrates a novel loss function specifically designed to simultaneously optimize fairness and model performance.
  2. 2.
    Based on FedFair, we propose an algorithm called FedFDP to further equip the system with differential privacy. In particular, we designed a fairness-aware gradient clipping strategy that ensures differential privacy and allows for adjustment of fairness in federated learning. In addition, we propose an adaptive clipping method for the additional loss values uploaded by each client to achieve optimal utility.
  3. 3.
    In addition, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
  4. 4.
    Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness. What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
II Related Work
II-A Differentially Private Federated Learning
Differentially Private Federated Learning (DPFL) has become widely used to enhance privacy protection in federated learning. Enayat et al. [45] dynamically adjusted compression rates in response to training loss changes to minimize communication rounds. Zhou et al. [52] examined the impact of local iterations when using the Laplace mechanism for perturbation. Fu et al. [13] introduced adaptive noise addition, customizing it based on loss fluctuations across clients and communication rounds. Ling et al. [31] balanced communication resources and privacy budgets by identifying the locally optimal number of iterations through convergence analysis. Raouf et al. [20] improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
II-B Fairness in DPFL
The above work in DPFL mainly focuses on model performance and privacy budget efficiency, without addressing fairness issues. In machine learning, a range of fairness concepts has been proposed, each emphasizing different aspects to meet diverse stakeholder needs. For instance, Good-Intent Fairness [38] and Selection Fairness [51] have been proposed. For a deeper understanding of these concepts, please refer to the relevant literature [7, 32, 34, 41].
In contrast, some recent studies have begun exploring group fairness within DPFL, aiming to reduce biases related to sensitive attributes such as gender and race [15, 39, 16, 17, 40]. Research suggests that adjusting parameters like clipping thresholds and noise multipliers can enhance group fairness in DPFL algorithms [16]. Under the framework of DP, adjacent datasets are defined as datasets $D$ and $D^{\prime}$ that differ by a single element, ensuring that an adversary cannot determine whether the output originates from $D$ or $D^{\prime}$. This mechanism significantly enhances the protection of attribute privacy and effectively reduces potential discrimination risks. Jagielski et al. [18] pioneered research on fair learning under DP constraints, addressing the need for predictive models to avoid discrimination based on specific attributes. They designed two learning algorithms that ensure both DP and equal opportunity, addressing challenges related to the collection of sensitive attributes. Tran et al. [44] utilized DP and Lagrangian duality to design neural networks capable of addressing fairness constraints while protecting the privacy of sensitive attributes. This approach ensures compliance with legal and ethical neutrality requirements when handling data related to gender, race, or age. Similarly, Ding et al. [9] proposed a classification model that combines a functional mechanism with decision boundary fairness to achieve both fairness and DP. This approach addresses privacy concerns related to training data and mitigates unfair practices stemming from attributes such as gender and race.
II-C Balanced Performance Fairness
Unlike the studies mentioned above, our research shifts focus to a relatively unexplored area within DPFL—balanced performance fairness. While group fairness emphasizes reducing performance biases related to sensitive attributes across different data types and is primarily used in centralized machine learning, balanced performance fairness, introduced by Li et al. [28], is more suited to federated learning. It emphasizes balancing the performance of global model across all clients, which is especially meaningful for data providers like companies and institutions. Building on the work of [28], Zhao et al. [50] proposed a strategy that replaces loss amplification with direct weight reallocation. Meanwhile, Li et al. [26] concentrated on improving fairness in personalized federated learning for local models.
Incorporating DP into model updates, with parameter adjustments such as clipping bounds and noise multipliers, can influence group fairness to some extent [16]. However, achieving balanced performance fairness necessitates adjustments to the direction of model updates. Therefore, determining how to update the model under DP constraints to ensure balanced performance fairness, while accounting for the effects of clipping and noise addition, presents a significant challenge. To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.
III Preliminaries
In the beginning, Table I presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
TABLE I: Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
III-A Federated Learning and Fairness
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$. The weights are defined as $p_{i}=\frac{|D_{i}|}{|D|}$, $\sum_{i=1}^{N}p_{i}=1$, $D_{i}$ represents the dataset of client $i$, and $D$ represents the union of all $D_{i}$.
We have defined balanced performance fairness in FL as follows.
Definition 1
(Balanced performance fairness [28]) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (2) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
III-B Differential Privacy
Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
Definition 2
(Differential Privacy [11]). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow. We use $l_{2}$ sensitivity as in the previous work.
DPSGD. In gradient-based methods, DPSGD[1] is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient for data sample $\xi_{j}$: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
  • Secondly, clipping the gradient of per-sample to get the sensitivity:
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$ (4)
    where $C$ represents the clipping bound.
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
The final step is to proceed with gradient descent as usual. When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
IV FedFair: Equipping FL with Fairness
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, fairness hyperparameter $\lambda$, batch sample ratio $q$, local dataset $D_{i}$.
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
5                   $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
13       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
14
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
20
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 1 FedFair
We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1. Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
Upon incorporating Equation (2) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
For client $i$ during round $t$, the refined objective is articulated as follows:
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (7)
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in Section VI-B. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (7), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (7) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
Figure 2: FedFDP framework.
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (IV) as show:
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Therefore, our approach can be recognized as establishing an dynamic learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as describe in Algorithm 1:
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round broadcasting the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
V FedFDP: Further Equipping FedFair with Differential Privacy
In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy. The flow of FedFDP algorithm is shown in Figure 2. It adds two additional processes to FedFair:
  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
Next, we discuss the details of these two strategies.
V-A Fair-clipping Strategy for Gradient
In the general DPSGD algorithm [1, 14, 4, 49], per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (7), into Equation (10):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (10)
where $\xi_{j}\in\mathcal{B}_{t}^{i}$, $j=\{1,2,\cdots,|\mathcal{B}_{i}|\}$, and $|\mathcal{B}_{i}|$ is the batch size of client $i$.
Next, we compute the gradient of Equation (10):
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (11)
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To ensure the algorithm is safeguarded by differential privacy, we replace the gradient of the loss function with Equation (11) and reformulate the gradient descent computation for client $i$ as follows:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta\tilde{g}^{i}_{t}$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
$=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (12)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (13)
Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (13) corresponds to the traditional DP clipping method without fairness. For large $C$, Equation (13) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (13) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
V-B Adaptive Clipping Method for Loss
Since calculating $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, differential privacy must be incorporated to ensure the algorithm preserves privacy. As shown in lines 2-2 of Algorithm 2, $F_{i}(\mathbf{w}_{t}^{i})$ must be clipped to the interval $[0,C_{l}^{i,t}]$, followed by the addition of Gaussian noise with a mean of 0 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, determining an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ prevents distortion of the original values but results in larger injected noise.
Based on the data heterogeneity of federated learning, we propose an adaptive clipping method for $F_{i}(\mathbf{w}_{t}^{i})$. For the $i$-th client, the adaptive clipping method uses the differentially private mean of the previous round as the clipping bound for the current round. The clipping bound $C_{l}^{i,t}$ for round $t$ and client $i$ is defined as follows, when $F_{i}(\mathbf{w}_{t-1}^{i})$ is the individual loss of the previous round and $\sigma^{t-1}_{l}$ is the individual noise multiplier of the previous round.
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (14)
where
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (15)
Input: loss function $F(\mathbf{w})$, learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$
1 SERVER
2       for $t\in[0,1,\cdots,T-1]$  do
3             for $i\in[1,2,\cdots,N]$ parallel  do
4                   $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\text{LOCAL}\left(\mathbf{w}_{t},F(\mathbf{w}_{t})\right)$;
5
6            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
7             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
8
9      return $\mathbf{w}_{T}$;
10
11LOCAL
12       Download $\text{Download }\mathbf{w}_{t},F(\mathbf{w}_{t})\text{ and }\mathbf{w}_{t}^{i}\leftarrow\mathbf{w}_{t}$;
13       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
14       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
15             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
16             Compute $C^{i,j}_{t}$ by Equation (13);
17             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
18
19      $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\tilde{g}_{t}^{i,j}+\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right)$;
20
21      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (14) and (15) ;
22       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
23             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
24
25            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
26
27      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
28       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
29
Algorithm 2 FedFDP
To ensure differential privacy during the clipping process, we use the noise-added loss $C_{l}^{i,t}$ as the clipping bound. For each communication round, the clipping bound follows the post-processing property of differential privacy.
V-C Overall Algorithm
We utilize the equation form Equation (12) to enhance Algorithm 2. Compared with the Algorithm 1, FedFDP solely necessitates the application of differential privacy to gradient and loss.
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
  • DP for local loss (lines 2-2): As described in Section V-B, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
VI Finding Optimal Fairness Parameter
In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
VI-A Convergence Analysis
We analyze the convergence of Algorithm 2 and derive insights on selecting the hyperparameter $\lambda$ based on the convergence upper bound. Prior to proving this, we need to establish several assumptions:
Assumption 1
$F_{1},\cdots,F_{N}$ are all L-smooth: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 2
$F_{1},\cdots,F_{N}$ are all $\mu$-strongly convex: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\geq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{\mu}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.
Assumption 3
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i\in[1,\cdots,N]$ , $t\in[0,\cdots,T-1]$ and $j\in[1,\cdots,|\mathcal{B}_{i}|]$
Theorem 1
Under the assumptions mentioned above, we obtain the convergence upper bound for Algorithm 2 as follows:
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (16)
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
The meaning of $\Gamma$ is consistent with [29], where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (16). For the detailed proof, please refer to Appendix A-A.
VI-B Optimal Fairness Parameter $\lambda^{*}$
Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (16) to obtain a function in terms of $\lambda$,
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (17)
where $a_{1}=G^{2}Q_{1}^{3}$, $a_{2}=6G^{2}Q_{1}^{2}$, $a_{3}=9Q_{1}G^{2}+2L\Gamma Q_{1}+2Q_{0}\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{4}=4G^{2}+2L\Gamma+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}$, $a_{5}=2Q_{0}$.
Let $\mathcal{F}(\lambda)=\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1}$, then $P(\lambda)=\frac{L}{2\mu T}\mathcal{F}(\lambda)$, where $\frac{L}{2\mu T}$ can be considered a constant term. Obviously, when $\mathcal{F}(\lambda)$ is minimized, $P(\lambda)$ is also minimized. The derivative of $\mathcal{F}(\lambda)$ is as follows:
$\mathcal{F}^{\prime}(\lambda)=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$ (18)
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
$\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$ (19)
Since $(a_{5}\lambda+1)^{2}>0$, the sign of $\mathcal{F}^{\prime}(\lambda)$ is consistent with that of $\mathcal{G}(\lambda)$.
Using the discriminant of a cubic function $\Delta=b^{2}-3ac$ (a general cubic function is expressed as $a\lambda^{3}+b\lambda^{2}+c\lambda+d$),
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
$=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$ (20)
The derivative of $\mathcal{G}(\lambda)$ is given by
$\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$ (21)
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
If $\mathcal{F}(\lambda)$ has an extremum, then $\mathcal{F}^{\prime}(\lambda)$ has a root, meaning $\mathcal{G}(\lambda)$ has a root. Substituting $\lambda=0$, $\mathcal{G}(0)=a_{3}-a_{4}a_{5}$, evidently when $a_{3}<a_{4}a_{5}$, $\mathcal{G}(\lambda)$ has a root.
Let $\lambda=x-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, replacing $\lambda$ with $x$ in $\mathcal{G}(\lambda)$ yields
$\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$ (22)
where
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
  • $a_{7}=\frac{108a_{1}^{2}a_{5}^{2}(a_{3}-a_{4}a_{5})-36(a_{1}a_{2}a_{5})(3a_{1}+a_{2}a_{5})+2(3a_{1}+a_{2}a_{5})^{3}}{216a_{1}^{3}a_{5}^{3}}$.
Using the root-finding formula to calculate $\mathcal{G}(x)=0$ gives the solutions $x_{1},x_{2},x_{3}$.
[…]
where imaginary number $\omega=\frac{-1+\sqrt{3}\mathbf{i}}{2}$, $\lambda_{l}=x_{l}-\frac{a_{2}a_{5}+3a_{1}}{6a_{1}a_{5}}$, $l=1,2,3$.
Furthermore, since $\lambda_{1}\lambda_{2}\lambda_{3}=-\frac{a_{3}-a_{4}a_{5}}{2a_{1}a_{5}}>0$, $\lambda$ has only one positive real root. If the equation has $k$ real roots, the optimal solution for $\lambda$ is $\lambda^{*}=\max_{k}(\lambda_{k})$.
In conclusion, $\mathcal{F}^{\prime}(\lambda)\geq 0\iff G(\lambda)\geq 0\iff\lambda\geq\lambda^{*}$, therefore $\mathcal{F}(\lambda)$ has a minimum value $\mathcal{F}(\lambda^{*})$, and $P(\lambda)$ has a minimum value $P(\lambda^{*})$.
VII Privacy Analysis
Since each client incurs the same privacy loss in every round, we evaluate privacy loss from the perspective of an individual client. Accordingly, we analyze the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 2
(Privacy loss of FedFDP). The privacy loss in $i$-th client of FedFDP satisfies:
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 3, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 4.
VII-A Proof of Theorem 2
Lemma 1
(Conversion from RDP to DP [3]). if a randomized mechanism $f:D\rightarrow\mathbb{R}$ satisfies $(\alpha,R)$-RDP ,then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
As Algorithm 2 shown, the $i$-th client uploads model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. Both $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the private training data of the client, and thus both require differential privacy protection, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
VII-A1 Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 3
After $T$ rounds local updates, the RDP of the $\mathbf{w}_{T}^{i}$ in $i$-th client satisfies:
$R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$ (24)
where $\sigma$ is noise multiplier of the $w_{t+1}^{i}$, and $\alpha>1$ is the order.
$proof.$ We will prove this in the following two steps: (i) use the RDP of the sampling Gaussian mechanism to calculate the privacy cost of each model update, and (ii) use the composition of RDP mechanisms to compute the privacy cost of multiple model updates.
Definition 3
(RDP privacy budget of SGM[37]). Let $SG_{q,\sigma}$, be the Sampled Gaussian Mechanism for some function $f$. If $f$ has sensitivity 1, $SG_{q,\sigma}$ satisfies $(\alpha,R)$-RDP whenever
$R\leq\frac{1}{\alpha-1}\log\max(A_{\alpha}(q,\sigma),B_{\alpha}(q,\sigma))$ (25)
where
$\left\{\begin{array}[]{l}A_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta_{0}}\left[\left(\vartheta(z)/\vartheta_{0}(z)\right)^{\alpha}\right]\\ B_{\alpha}(q,\sigma)\triangleq\mathbb{E}_{z\sim\vartheta}\left[\left(\vartheta_{0}(z)/\vartheta(z)\right)^{\alpha}\right]\end{array}\right.$ (26)
with $\vartheta_{0}\triangleq\mathcal{N}\left(0,\sigma^{2}\right),\vartheta_{1}\triangleq\mathcal{N}\left(1,\sigma^{2}\right)\mbox{ and }\vartheta\triangleq(1-q)\vartheta_{0}+q\vartheta_{1}$
Further, it holds $\forall(q,\sigma)\in(0,1],\mathbb{R}^{+*},A_{\alpha}(q,\sigma)\geq B_{\alpha}(q,\sigma)$. Thus, $SG_{q,\sigma}$ satisfies $\left(\alpha,\frac{1}{\alpha-1}\log\left(A_{\alpha}(q,\sigma)\right)\right)$-RDP .
Finally, the existing work [37] describes a procedure to compute $A_{\alpha}(q,\sigma)$ depending on integer $\alpha$.
$A_{\alpha}=\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right)$ (27)
Definition 4
(Composition of RDP[36]). For two randomized mechanisms $f,g$ such that $f$ is $(\alpha,R_{1})$-RDP and $g$ is $(\alpha,R_{2})$-RDP the composition of $f$ and $g$ which is defined as $(X,Y)$(a sequence of results), where $X\sim f$ and $Y\sim g$, satisfies $(\alpha,R_{1}+R_{2})-RDP$
From Definition 3 and Definition 4, the Theorem 3 is obtained.
VII-A2 Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$.
Theorem 4
After $T$ rounds local updates, the RDP of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ in $i$-th client satisfies:
$R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-k}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$ (28)
where $\sigma_{l}$ is noise multiplier of the $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$, and $\alpha>1$ is the order.
The proof is similar to that of the $w_{t+1}^{i}$, so we omit it.
VII-A3 Proof Loss of FedFDP.
Since both $w_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 4, and then use Lemma 1 to convert it to $(\epsilon,\delta)$-DP. Therefore, the Theorem 2 is hold.
VIII Experiments
TABLE II: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
FedAvg 98.67$\pm$0.09 7.3e-3$\pm$5.4e-3 87.05$\pm$1.02 3.6e-1$\pm$2.2e-1 62.03$\pm$1.00 1.1e0$\pm$8.9e-1
SCAFFLOD 98.45$\pm$0.03 3.0e-2$\pm$2.6e-2 86.77$\pm$0.25 6.1e0$\pm$4.0e0 62.21$\pm$1.37 9.9e0$\pm$1.1e1
FedProx 98.70$\pm$0.13 7.5e-3$\pm$5.1e-3 87.30$\pm$0.51 3.5e-1$\pm$2.1e-1 62.13$\pm$0.20 1.2e0$\pm$3.1e-1
FedDyn 98.40$\pm$0.23 8.0e-3$\pm$2.1e-3 87.48$\pm$0.31 6.2e-1$\pm$3.5e-1 62.33$\pm$1.15 3.2e0$\pm$8.8e-1
q-FFL 98.72$\pm$0.03 4.7e-3$\pm$1.0e-3 87.35$\pm$0.13 5.1e-1$\pm$5.2e-2 63.33$\pm$0.21 9.4e-1$\pm$3.3e-2
FedFair 98.75$\pm$0.09 3.9e-3$\pm$1.3e-3 87.70$\pm$0.76 2.6e-1$\pm$2.3e-1 62.38$\pm$0.88 8.5e-1$\pm$6.6e-1
TABLE III: FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
Heterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
FedAvg 98.67 7.3e-3 98.37 1.2e-3 89.41 3.6e-2 62.03 1.1e0 60.89 1.5e0 59.78 1.5e0
SCAFFLOD 98.45 5.0e-3 98.51 2.1e-3 86.39 3.4e-2 62.21 9.9e0 59.48 8.7e0 58.64 2.3e0
FedProx 98.70 7.5e-3 98.35 1.2e-3 88.33 3.5e-2 62.13 1.2e0 60.33 2.2e0 61.31 1.6e0
FedDyn 98.40 8.0e-3 98.36 1.8e-3 88.26 3.5e-2 62.33 9.4e-1 61.35 2.6e0 62.52 1.5e0
q-FFL 98.72 4.7e-3 98.61 1.1e-3 89.03 3.6e-2 63.33 9.4e-1 60.63 1.3e0 59.65 1.5e0
FedFair 98.75 3.9e-3 98.72 1.2e-3 89.60 1.7e-2 62.38 8.5e-1 61.58 1.1e0 62.90 1.5e0
In this section, we present the comparative experimental results of FedFair and FedFDP with multiple baseline methods on three datasets: MNIST [22], FashionMNIST [48], and CIFAR10 [21]. We provide a detailed description of the experimental evaluation of FedFair and FedFDP in terms of performance, fairness, heterogeneity, and scalability. In particular, we conduct ablation studies on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$, and noise multiplier $\sigma$. In all tables, $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are better.
We implement our experiment using PyTorch-1.8 and run all experiments on a server with one Intel i9 13900ks CPU (24 cores), 64GB of memory, and one NVIDIA 4090 GPU, running Windows 10.
VIII-A Default Settings
We compared our algorithms against six baseline methods. For FedFair, the baseline algorithms were left unmodified; for FedFDP, DP was added to each baseline algorithm to enable a fair comparison.
  1. 1.
    FedAvg[33]: FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. 2.
    SCAFFLOD[19]: SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. 3.
    FedProx[27]: FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. 4.
    FedDyn[2]: FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. 5.
    ALI-DPFL[31]: ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. 6.
    q-FFL[28]: q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings [30, 24] to 10 clients which controlled by a Dirichlet distribution denoted as $\mathrm{Dir}(\beta)$, and the default value of $\beta=0.1$ [30, 46]. We use a 4-layer CNN architecture [33] which consists of two convolutional layers and two fully connected layers as model architecture.
Hyperparameters. For FedFair and its baseline algorithms, we set learning rate $\eta=0.1$. For FedFDP and its baseline algorithms, we set the $\eta=1.0$, the batch sample ratio $q=0.05$, clipping bound for gradient $C=0.1$, the noise multipliers for gradient $\sigma=2.0$, the privacy budget $\epsilon=3.52$ and the $\delta=1e-5$. In particular, for the loss values in the FedFDP algorithm that additionally require differential privacy processing, we set the clipping bound for loss $C_{l}=2.5$ and the noise multipliers for loss $\sigma_{l}=5.0$.
TABLE IV: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
Datasets MNIST FashionMNIST CIFAR10
Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
FedAvg 93.40$\pm$0.47 1.1e11$\pm$0.5e10 84.15$\pm$1.97 3.0e8$\pm$1.4e8 52.61$\pm$0.17 6.1e9$\pm$5.2e9
SCAFFLOD 93.95$\pm$1.39 7.0e10$\pm$4.9e9 83.41$\pm$4.87 3.3e9$\pm$4.7e9 53.85$\pm$1.16 6.2e9$\pm$1.3e9
FedProx 90.50$\pm$3.95 5.3e10$\pm$7.3e9 83.85$\pm$0.98 7.9e8$\pm$4.5e8 51.34$\pm$0.91 4.4e9$\pm$2.1e9
FedDyn 91.42$\pm$2.25 5.6e10$\pm$9.8e9 84.04$\pm$1.21 8.8e8$\pm$3.2e8 52.14$\pm$1.39 5.5e9$\pm$1.2e9
ALI-DPFL 90.89$\pm$1.65 3.3e10$\pm$8.3e9 83.65$\pm$1.68 6.6e8$\pm$3.5e8 52.05$\pm$1.26 3.6e9$\pm$1.6e9
q-FFL 93.74$\pm$1.41 7.8e10$\pm$1.1e11 83.13$\pm$2.74 4.2e9$\pm$2.6e9 48.46$\pm$1.00 4.7e9$\pm$1.9e9
FedFDP 95.13$\pm$0.83 2.3e10$\pm$1.0e10 85.99$\pm$0.76 2.8e8$\pm$0.8e8 54.21$\pm$0.98 2.6e9$\pm$1.6e9
TABLE V: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
FedAvg 93.40 1.1e11 94.17 4.1e9 85.60 3.7e9 52.61 6.1e9 51.58 2.0e9 55.90 1.5e9
SCAFFLOD 93.95 7.0e10 94.06 1.9e9 84.56 3.6e9 53.85 6.2e9 54.34 2.9e8 53.98 2.2e9
FedProx 90.50 5.3e10 94.23 3.1e9 84.38 5.6e9 51.34 4.4e9 53.67 4.5e8 54.14 2.2e9
FedDyn 91.42 5.6e10 93.08 3.8e9 85.21 9.8e10 52.14 5.5e9 50.21 1.2e10 51.86 2.0e10
ALI-DPFL 90.89 3.3e10 93.98 2.6e9 83.66 2.5e10 52.05 3.6e9 38.35 2.4e10 33.46 1.6e10
q-FFL 93.74 7.8e10 93.86 2.5e8 85.20 7.4e8 48.46 4.7e9 51.12 2.7e8 49.83 1.1e7
FedFDP 94.13 2.3e10 94.56 3.4e9 86.77 6.0e8 54.21 2.6e9 54.52 2.2e8 56.49 5.3e8
[…]
$\epsilon$ 1.0 2.0 3.0 4.0
Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$ Acc.($\uparrow$) $\Psi(\downarrow)$
FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
VIII-B Result of FedFair
Performance and Fairness. Table II presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%. A plausible explanation is that, as the training process progresses, FedFair increasingly prioritizes fairness, resulting in a reduction in $\eta_{i}$. Compared to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods [55].
Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
Figure 4: Impact of different $C$ in FedFDP algorithm.
Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
Heterogeneity. Table III shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table III indicates that FedFair generally surpasses baselines in accuracy and shows fairness improvements of 9.6%, 15.4%, and 0.01% across three client counts.
VIII-C Results of FedFDP
Performance and Fairness. Table IV displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table V shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table V demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Impact of Different $\lambda$. We investigated the correlation between $\lambda$ and $\Psi$ in the FedFDP algorithm. As depicted in Figure 3, our findings align with the theoretical analysis: an initial increase in $\lambda$ leads to a decrease in $\Psi$, followed by an increase, indicating the existence of an optimal $\lambda$.
Impact of Different $C$. We emphasized that excessively large or small values of $C$ cannot ensure both performance and fairness simultaneously. As illustrated in Figure 4, we examined the correlation between $C$, performance, and fairness across two datasets, with $C$ taking values in the range [0.01, 0.1, 1, 5, 10]. Although $\Psi$ reaches its minimum when $C=1$, the model does not converge due to excessive noise addition. As $C$ continues to increase, the model’s performance further deteriorates, resulting in an increase in $\Psi$. When $C=0.1$, both accuracy and $\Psi$ achieve relatively good performance.
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table VI. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
Impact of Different $\sigma$. We tested the variations in accuracy and fairness on two datasets when $\sigma$ was set to $[1,1.5,2,2.5,3]$, under the condition of $\epsilon=2$ and $q=0.05$, which supports $T=[6,115,268,463,708]$ rounds of communication. As shown in Figure 5, an increase in $\sigma$ allows more training epochs to advance the model’s convergence process, but when $\sigma=3$, excessive noise hinders the convergence. When $\sigma$ is smaller, the model performance is relatively similar, with a lower $\Psi$ value. As $\sigma$ gradually increases, the performance differences between models become apparent, and the $\Psi$ value increases.
IX Conclusion
In this paper, a FedFair algorithm is initially proposed to effectively address fairness concerns by optimizing a novel local loss function. The FedFDP, building upon FedFair, is introduced to incorporate a fairness-aware gradient clipping strategy and an adaptive clipping method for additional loss values, thereby achieving both fairness and differential privacy protection. Then, we find an optimal fairness parameter $\lambda^{*}$ through convergence analysis and numerical analysis methods, striking a balance between model performance and fairness. Subsequently, a comprehensive privacy analysis of the approach is conducted using RDP. Through extensive experiments, the results indicate that FedFair and FedFDP significantly outperform state-of-the-art solutions in terms of model performance and fairness. It is believed that our work contributes a valuable FL framework for addressing fairness and privacy challenges.
Appendix A Appendix
A-A Proof of Theorem 1
For the purpose of validation, we introduce an additional variable $\mathbf{v}_{t}^{i}$ to represent the immediate result of a single-step DPSGD update from $\mathbf{w}_{t}^{i}$. We interpret $\mathbf{w}_{t+1}^{i}$ as the parameter obtained after a single communication step. Consequently, the fair-clipping DPSGD in client $i$ at iteration $t$ transitions from Equation (12) to:
$\mathbf{v}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (29)
where:
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
In our analysis, we define two virtual sequences $\mathbf{v}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{v}_{t}^{i}$ and $\mathbf{w}_{t}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t}^{i}$, which is motivated by [43]. Therefore,
$\mathbf{v}_{t+1}=\mathbf{w}_{t}\quad-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$ (30)
A-A1 Key Lemma
Lemma 2
(Results of one iteration.)Assume Assumption 1-3 hold, we have:
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq\left(1-\mu\eta C_{t}\right)\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where:
  • $A\triangleq G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}$,
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
  • $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i},C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Let $\Delta_{t}\triangleq\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. It is evident that we always have $\mathbf{w}_{t+1}=\mathbf{v}_{t+1}$. According to Lemma 2, this implies:
$\Delta_{t+1}\leq\left(1-\mu\eta C_{t}\right)\Delta_{t}+\eta^{2}A$
We use mathematical induction to obtain $\Delta_{t}\leq\frac{v}{t}$ where $v=\max\{\frac{\beta^{2}A}{\mu\beta C_{t}-1},\Delta_{1}\}$, $\eta=\frac{\beta}{t}$ for some $\beta>\frac{1}{\mu}$.
STEP 1. When $t=1$, the equation $\Delta_{1}\leq v$ holds obviously.
STEP 2. We assume $\Delta_{t}\leq\frac{v}{t}$ holds.
STEP 3.
$\Delta_{t+1}\leq\left(1-\mu\frac{\beta}{t}C_{t}\right)\frac{v}{t}+\frac{\beta^{2}A}{t^{2}}=\frac{t-1}{t^{2}}v+\left(\frac{\beta^{2}A}{t^{2}}-\frac{\mu\beta C_{t}-1}{t^{2}}v\right)\leq\frac{t-1}{t^{2}}v\leq\frac{v}{t+1}$
Therefore, $\Delta_{t+1}\leq\frac{v}{t+1}$ holds, completing the proof by mathematical induction. Hence, $\Delta_{t}\leq\frac{v}{t}$ holds.
Then by the $L$-smoothness of $F(\cdot)$, let $\beta=\frac{2}{\mu}$, we get
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2}\Delta_{t}\leq\frac{L}{2t}v\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\Delta_{1}\right)$
A-A2 Proof of Lemma 2
By the Equation (30), we get
$\left\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\right\|^{2}$
$=\|\mathbf{w}_{t}-\mathbf{w}^{*}-\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\|^{2}$
$=\left\|\mathbf{w}_{t}-\mathbf{w}^{*}\right\|^{2}$
$\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\mathcal{N}(0,\mathbf{I})]}_{\mathcal{A}_{1}}\rangle$
$+\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}}_{\mathcal{A}_{2}}$
Firstly, we process $\mathcal{A}_{2}$:
$\mathcal{A}_{2}\leq\underbrace{\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2}}_{\mathcal{B}_{1}}$
$\quad+\underbrace{\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\left\langle\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),\sigma C\cdot\mathcal{N}(0,\mathbf{I})\right\rangle}_{\mathcal{B}_{0}}$
$\quad+\underbrace{\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\cdot\mathcal{N}(0,\mathbf{I})\|^{2}}_{\mathcal{B}_{2}}$
Since $\mathbb{E}[\mathcal{B}_{0}]=0$, we focus on $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$:
$\mathbb{E}[\mathcal{B}_{2}]\leq\frac{\eta^{2}}{\hat{B}^{2}}\sum_{i=1}^{N}\mathbb{E}\|\sigma C\mathcal{N}(0,\mathbf{I})\|^{2}\leq\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}},$
where $\frac{1}{\hat{B}^{2}}\triangleq\max_{i}\frac{1}{|\mathcal{B}_{i}|}$.
By the convexity of $\|\cdot\|^{2}$,
$\mathcal{B}_{1}\leq\eta^{2}\sum_{i=1}^{N}p_{i}\left\|\frac{1}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\right\|^{2},$
taking exception and according to assumption 3:
$\mathbb{E}[\mathcal{B}_{1}]\leq\eta^{2}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\left\|C_{t}^{i}\nabla F_{i}(\mathbf{w}_{t}^{i})\right\|^{2}\right]\leq\eta^{2}C_{t}^{2}G^{2},$
where $C_{t}\triangleq\sum_{i=1}^{N}p_{i}C_{t}^{i}$, $C_{t}^{i}\triangleq\frac{1}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}$.
Now, we obtain the bound for the expectation of $\mathcal{A}_{2}$:
$\mathbb{E}[\mathcal{A}_{2}]$ $\leq\eta^{2}\left(\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}+G^{2}C_{t}^{2}\right)$
The process of $\mathcal{A}_{1}$ show as below:
$\mathcal{A}_{1}$ $=-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle\underbrace{-2\langle\mathbf{w}_{t}-\mathbf{w}^{*},\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sigma C\mathcal{N}(0,\mathbf{I})\rangle}_{\mathcal{C}_{0}}$
$=\mathcal{C}_{0}\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{1}}\underbrace{-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle}_{\mathcal{C}_{2}}$
It’s obvious that $\mathbb{E}[\mathcal{C}_{0}]=0$.
By Cauchy-Schwarz inequality and AM-GM inequality, we have
$\mathcal{C}_{1}$ $=-2\sum_{i=1}^{N}p_{i}\langle\mathbf{w}_{t}-\mathbf{w}_{t}^{i},\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\rangle$
$\leq\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|^{2}$
$=\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\mathcal{B}_{1}$
So we get
$\mathbb{E}[\mathcal{C}_{1}]\leq\sum_{i=1}^{N}p_{i}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}$
According to Assumption 2, we know that
$-\langle\mathbf{w}_{t}^{i}-\mathbf{w}^{*},\nabla F_{i}(\mathbf{w}_{t}^{i})\rangle\leq-\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
So we get
$\mathcal{C}_{2}$ $\leq 2\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}\sum_{j=1}^{|\mathcal{B}_{i}|}C_{t}^{i,j}\cdot\left[-\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F_{i}(\mathbf{w}^{*})\right)-\frac{\mu}{2}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}\right]$
$\leq-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}+F^{*}-F_{i}(\mathbf{w}^{*})\right)-\mu\eta C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}^{*}\|^{2}$
$=-2\eta C_{t}\underbrace{\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F^{*}\right)}_{\mathcal{D}_{1}}-2\eta C_{t}\Gamma-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2},$
where $\Gamma\triangleq\sum_{i=1}^{N}p_{i}(F^{*}-F_{i}^{*})=F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$.
Next, we proceed to handle $\mathcal{D}_{1}$.
$\mathcal{D}_{1}$ $=\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}(\mathbf{w}_{t})\right)+\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w}_{t})-F^{*}\right)$
$\geq\sum_{i=1}^{N}p_{i}\langle\nabla F_{i}(\mathbf{w}_{t}),\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\rangle+(F(\mathbf{w}_{t})-F^{*})$
(from the Assumption 2)
$\geq-\frac{1}{2}\sum_{i=1}^{N}p_{i}\left[\eta\|\nabla F_{i}(\mathbf{w}_{t})\|^{2}+\frac{1}{\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]+(F(\mathbf{w}_{t})-F^{*})$
(from the AM-GM inequality)
$\geq-\sum_{i=1}^{N}p_{i}\left[\eta L(F_{i}(\mathbf{w}_{t})-F_{i}^{*})+\frac{1}{2\eta}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]+(F(\mathbf{w}_{t})-F^{*})$
(from the L-smooth inference)
$\geq-(\eta L+1)\Gamma-\frac{1}{2\eta}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2},$
where L-smooth inference as show:
$\|\nabla F_{i}(\mathbf{w}_{t}^{i})\|^{2}\leq 2L\left(F_{i}(\mathbf{w}_{t}^{i})-F_{i}^{*}\right).$ (31)
Thus, we get
$\mathcal{C}_{2}\leq 2\eta^{2}C_{t}L\Gamma+C_{t}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}-\mu\eta C_{t}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
To sum up,
$\mathbb{E}[\mathcal{A}_{1}]$ $=\mathbb{E}\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}+\eta^{2}G^{2}C_{t}^{2}+2\eta^{2}C_{t}L\Gamma$
$\quad+C_{t}\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{w}_{t}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$=(1+C_{t})\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]-\mu\eta C_{t}\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$
$\quad+\eta^{2}\left(G^{2}C_{t}^{2}+2L\Gamma C_{t}\right),$
and
$\quad\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}-\mathbf{w}_{t}^{i}\|^{2}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|(\mathbf{w}_{t}-\mathbf{v}_{t+1}^{i})-(\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i})\|^{2}\right]$
$\leq\mathbb{E}\left[\sum_{i=1}^{N}p_{i}\|\mathbf{w}_{t}^{i}-\mathbf{v}_{t+1}^{i}\|^{2}\right]$
$\leq\mathbb{E}\left\|\sum_{i=1}^{N}p_{i}\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]\right\|^{2}$
$=\mathbb{E}\left[\mathcal{A}_{2}\right]=\eta^{2}C_{t}^{2}G^{2}+\frac{\eta^{2}\sigma^{2}C^{2}d}{\hat{B}^{2}}.$
So we get,
$\mathbb{E}\left[\mathcal{A}_{1}\right]$ $=(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+(1+C_{t})\eta^{2}C_{t}^{2}G^{2}+\eta^{2}\left(2G^{2}C_{t}^{2}+2L\Gamma C_{t}+2\frac{\sigma^{2}C^{2}d}{\hat{B}^{2}}\right)$
All in all, we get
$\mathbb{E}\|\mathbf{v}_{t+1}-\mathbf{w}^{*}\|^{2}\leq(1-\mu\eta C_{t})\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}+\eta^{2}A,$
where
$A=G^{2}C_{t}^{3}+3G^{2}C_{t}^{2}+2L\Gamma C_{t}+\frac{2\sigma^{2}C^{2}d}{\hat{B}^{2}}.$